OpenStudy (anonymous):

X/2=3/6 Solve this proportion

OpenStudy (anonymous):

x/2=3/6 x/2=1/2

OpenStudy (anonymous):

-3/4=a/12

OpenStudy (anonymous):

I am sorry I have to go.

OpenStudy (anonymous):

\[-\frac{3}{4} = \frac{a}{12}\] What you want to do is make the first equation's denominator equal the 2nd one. What can you multiply by -4 to get 12?

OpenStudy (anonymous):

3 any tool for this too hahaha

OpenStudy (anonymous):

Right, so you want to multiply the top and the bottom by -3 on the first fraction. \[-\frac{3}{4}\times \frac{3}{3}\] Can you solve that? Hahah, not that I've come across. Even if there was, it wouldn't benefit you in the long run. Cheat the homework, fail the test :D

OpenStudy (anonymous):

I'm sorry, the -3 in my first sentence is supposed to be a positive 3. Oops

OpenStudy (anonymous):

9/12 3/4

OpenStudy (anonymous):

Right, but it'd be -9/12, right? So a = -9

OpenStudy (anonymous):

tht he answre

OpenStudy (anonymous):

That's the answer. Do you understand how we got that?

OpenStudy (anonymous):

yuh

OpenStudy (anonymous):

You told me that \[-\frac{3}{4} = -\frac{9}{12}\]\[\frac{-9}{12} = \frac{a}{12}\] \[a = -9\]